OpenStudy (anonymous):

Factor completely: 3x^2 + 2x - 1 (3x + 1)(x - 1) (3x + 1)(x + 1) (3x - 1)(x + 1) (3x - 1)(x - 1)

OpenStudy (anonymous):

its the 3rd option

OpenStudy (anonymous):

x² + 1/3x = 1/3 + (1/3)² x² + 1/3x = (3 + 1)/9 (x + 1/3)² = 4/9 x + 1/3 = ± 2//3 = x + 1/3 - 2/3, = x - 1/3, = 3x - 1 = x + 1/3 + 2/3, = x + 3/3, = x + 1

OpenStudy (anonymous):

Factor 3x^2+2x−1: 1. Multiply 3 by -1, which is -3. 2. Which two numbers add up to 2 and multiply to -3? 3. Answer: 3 and -1 4. Rewrite 2x as the sum of 3x and −x: 3x^2+3x−x−1 2 Factor out common terms in the first two terms, then in the last two terms 3x(x+1)−(x+1) 3 Factor out the common term x+1 (x+1)(3x−1)

OpenStudy (dannyrod2000):

(3x-1)(x+1)

OpenStudy (anonymous):

Thanks All!!